OpenStudy (anonymous):

solve each system of equations 4a-3b=17 5a+4b-16

OpenStudy (anonymous):

x ist equn. by 4 and second by 3: 16a - 12b = 68 15a + 12b = 48 add 31a = 116 a = 116/31 = 3.7419 plug this into first equation 4 * 3.7419 - 3b = 17 b = (4 * 3.7419 - 17) / 3 = -0.6774

OpenStudy (anonymous):

huh ?

OpenStudy (anonymous):

what does a EQUAL and what does B equal ?

OpenStudy (anonymous):

a = 3.7419 b = -0.6774

OpenStudy (anonymous):

do u understand how it was done? - ill willingly explain

OpenStudy (anonymous):

please do jiimmyrep

OpenStudy (anonymous):

you can add or subtract equations just like numbers or albegraic terms you do this to make the numbers in front of the 'a' or 'b' terms the same i chose the b terms -by multiplying the first equation by 4 and second by 3 i got -12b and +12b. Note you have to multiply every term in the equations. Now by adding the new equations the 'b' terms are eliminated and you get an equation in a only. solve this for a . Then plug the values of a into one of the first equations to get the value of b.